# Data-Efficient Molecular Generation with Hierarchical Textual Inversion

### v1

@@ Abstract
Developing an effective molecular generation framework even with a limited number of molecules is often important for its practical deployment, e.g., drug discovery, since acquiring task-related molecular data requires expensive and time-consuming experimental costs. To tackle this issue, we introduce Hierarchical textual Inversion for Molecular generation (HI-Mol), a novel data-efficient molecular generation method. HI-Mol is inspired by the importance of hierarchical information, e.g., both coarse- and fine-grained features, in understanding the molecule distribution. We propose to use multi-level embeddings to reflect such hierarchical features based on the adoption of the recent textual inversion technique in the visual domain, which achieves data-efficient image generation. Compared to the conventional textual inversion method in the image domain using a single-level token embedding, our multi-level token embeddings allow the model to effectively learn the underlying low-shot molecule distribution. We then generate molecules based on the interpolation of the multi-level token embeddings. Extensive experiments demonstrate the superiority of HI-Mol with notable data-efficiency. For instance, on QM9, HI-Mol outperforms the prior state-of-the-art method with 50$\times$ less training data. We also show the effectiveness of molecules generated by HI-Mol in low-shot molecular property prediction.
@@ 1 Introduction
Figure 1: Overview of our HI-Mol framework for data-efficient molecular generation. (1) Hierarchical textual inversion: we encode the features of molecules into multi-level token embeddings. (2) Embedding interpolation-based sampling: we generate novel molecules using interpolation of low-level token embeddings.
@@
We learn this hierarchical latent embedding in the latent space of the recent text-to-molecule model (Edwards et al., 2022). Specifically, we adopt textual inversion (Gal et al., 2022)—which learns low-shot image distribution by introducing a new single token within a text-to-image model into this text-to-molecule model with hierarchy (see Figure 1). A key difference here is that the consideration of “multi-level” tokens are essential in learning the low-shot molecule distribution due to the complicatedly structured nature of molecules compared to images (see Table 1).
@@ 2 Related Work
Molecular generation. Most molecular generation methods fall into three categories. First, graph-based methods (Hoogeboom et al., 2022; Zhang et al., 2023; Vignac et al., 2023) formalize molecular generation as a graph generation problem by representing each molecule as an attributed graph. Next, fragment-based methods (Jin et al., 2018; Kong et al., 2022) define a dictionary of chemically meaningful fragments, e.g., functional groups. A molecule is represented as a tree structure of dictionary elements and the distribution of connected fragments is then modeled. Finally, string-based methods (Gómez-Bombarelli et al., 2016; Flam-Shepherd et al., 2022; Ahn et al., 2022) utilize the Simplified Molecular-Input Line-Entry System (SMILES, Weininger, 1988) to write molecules as strings and learn the distribution of molecules in this string space. Our method takes the string-based approach based on the recent large-scale text-to-molecule models that use the SMILES representation.
Hierarchical generation methods. Recent molecular generation methods introduce the notion of hierarchy in molecular generation (Jin et al., 2020; Zhu et al., 2023). Specifically, they incorporate the hierarchy within a single molecule (e.g., atom- and motif-level). However, they do not consider the hierarchy among molecules (e.g., dataset-, cluster-, and molecule-level), which is indeed crucial to understand the molecular dataset, i.e., target distribution (Alexander et al., 2011). To overcome this limitation, we carefully design our multi-level embeddings to reflect the hierarchy among the target molecules through our hierarchical tokens.
@@
In contrast to the recent advances of low-shot generation in the image domain, developing a low-shot (or data-efficient) molecular generation framework is relatively under-explored despite its importance in practical applications (Altae-Tran et al., 2017; Guo et al., 2022). Hence, our method tackles this problem by designing a molecule-specific inversion method using the recent large-scale text-to-molecule models. Specifically, we incorporate the concept of “hierarchy” of molecular structures (Alexander et al., 2011) into the textual inversion framework. Due to this unique motivation, our method effectively learns the molecule distribution of low-shot molecules with diverse molecular structures, while the applications of prior works, e.g., Guo et al. (2022), are limited to structurally similar low-shot molecules such as monomers and chain-extenders.
@@ 3 HI-Mol: Hierarchical Textual Inversion for Molecular Generation
In Section 3.1, we provide an overview of our problem and the main idea. In Section 3.2, we provide descriptions of textual inversion to explain our method. In Section 3.3, we provide a component-wise description of our method.
@@ 4.2 Main Results
We also propose a simple algorithm to modify the generated invalid SMILES strings by correcting invalid patterns [^5] For example, we modify the invalid SMILES caused by the unclosed ring, e.g., $\mathtt{C1CCC}\rightarrow\mathtt{CCCCC}$. Please see Appendix J for the detailed algorithm. We mark in the Grammar column in Table 2 and 4 when modification is applied for evaluation. without a computational overhead. By applying this modification algorithm, we convert an invalid SMILES string to a valid one that represents a valid molecule, therefore, the Validity score becomes 100 in this case. In particular, the molecules from the modified SMILES further improve the overall metrics, e.g., FCD by 19.0 $\rightarrow$ 16.6 and 11.2 $\rightarrow$ 10.7 in the HIV and the BBBP dataset, respectively. This indicates that the modified SMILES indeed represent molecules from the desired low-shot molecule distribution and further highlights the superior quality of our generated molecules.
@@
For an extensive comparison with the baselines which show high Uniqueness and Novelty scores, e.g., GDSS (Jo et al., 2022), we perform an additional comparison after we adjust the Uniqueness and Novelty scores of our method to 100; this setup allows us to perform a fair comparison in the FCD score with these methods. Here, we adjust the sampling strategy slightly; we ignore the generated molecules which have an overlap with the training molecules and the already generated molecules. Even in this case, HI-Mol achieves an FCD of 0.601, which outperforms all these baselines. We provide detailed results and discussion in Appendix I.
@@ 4.3 Analysis
Ablation on hierarchical textual inversion. To validate the effectiveness of each component in our HI-Mol framework, we compare the results where some components are excluded from the overall framework. Specifically, we compare the generation performance of the following setups: (1) not using the inversion technique; we train the text-to-molecule model with the molecule-description pairs, (2) using the shared token $[S^{\ast}]$ only, (3) using $[S^{\ast}]$ and the detail tokens $[D_{n}^{\ast}]$, (4) using all three types of tokens, and (5) applying the additional modification algorithm. Note that for (1) and (2), it is impossible to apply our interpolation-based sampling; instead, we use temperature sampling with temperature $\tau=2.0$. We provide this result in Table 6. First, we find that (1) the naïve training and (2) the inversion with a single shared token (Gal et al., 2022) do not show reasonable performance, i.e., they achieve only 0.4% Validity. In (3) and (4), introducing low-level tokens in the inversion framework significantly improves the generation quality by learning the low-level features in molecules. Finally, (5) the modification algorithm converts an invalid generated SMILES into a valid one that lies in our target distribution.
@@ Appendix A Method Details
We utilize a recently introduced text-to-molecule model, MolT5-Large-Caption2Smiles (Edwards et al., 2022) in our HI-Mol framework. [^7] https://huggingface.co/laituan245/molt5-large-caption2smiles This model is constructed upon a text-to-text model, T5 (Raffel et al., 2020), and molecular information is injected by additional training with both unpaired SMILES (Weininger, 1988) string and caption-SMILES paired dataset. Our experiment is conducted for 1,000 epochs using a single NVIDIA GeForce RTX 3090 GPU with a batch size of 4. We use AdamW optimizer with $\epsilon=1.0\times 10^{-8}$ and let the learning rate $0.3$ with linear scheduler. We clip gradients with the maximum norm of 1.0. We update the assigned cluster $c_{n}$ of each molecule for the first 5 epochs following Eq. (3.3). For interpolation-based sampling, we choose a uniform distribution $p(\lambda)$, (i.e., $p(\lambda)\coloneqq\mathcal{U}(l,1-l)$), where $\lambda$ controls relative contributions of interpolated token embeddings. We set $l=0.0$ on the datasets in MoleculeNet benchmark (Wu et al., 2018), and $l=0.3$ on the QM9 dataset (Ramakrishnan et al., 2014).
@@ Appendix H Additional Experiments
Table 14: Quantitative results of the few-shot generation experiments on subsets of the HIV dataset (Wu et al., 2018). We generate the same number of molecules as the number of the training samples. Due to the large training cost, we report the score of DEG (Guo et al., 2022) only for 30 samples.
30 DEG (Guo et al., 2022) Graph 03.3 39.2 0.105 100 100 100
STGG (Ahn et al., 2022) SMILES 00.0 41.5 0.110 100 67 100
CRNN (Segler et al., 2018) SMILES 00.0 40.0 0.121 80 71 100
HI-Mol (Ours) SMILES 08.3 34.8 0.103 80 75 100
150 STGG (Ahn et al., 2022) SMILES 01.3 28.2 0.054 100 90 100
CRNN (Segler et al., 2018) SMILES 01.3 30.1 0.063 50 84 100
HI-Mol (Ours) SMILES 08.3 22.1 0.038 64 91 100
500 STGG (Ahn et al., 2022) SMILES 01.3 22.8 0.041 100 74 100
CRNN (Segler et al., 2018) SMILES 02.7 30.0 0.064 51 100 100
HI-Mol (Ours) SMILES 10.3 20.8 0.020 63 91 100
@@
Extremely limited data regime. Since our model exploits the power of large molecular language models by designing a molecule-specialized textual inversion scheme, one can expect our model to be beneficial in extremely limited data regimes compared with prior methods. To verify this, we conduct an experiment using only a subset of the HIV dataset and report its quantitative result in Table 14. Even with this situation, HI-Mol still outperforms prior state-of-the-art molecular generation methods, e.g., our method improves FCD as 39.2 $\to$ 34.8 when trained with 30 samples.
@@
Appendix I Details on QM9 Experiments
[…]
2 0.434 0.001 90.7 75.8 73.5
0.430 0.001 100 76.1 75.6
5 0.412 0.001 89.4 85.8 70.4
0.410 0.001 100 86.4 72.4
10 0.400 0.002 87.6 87.6 71.2
0.398 0.001 100 88.3 73.2
20 0.384 0.001 86.7 87.8 70.0
0.383 0.001 100 88.7 71.8
50 0.372 0.001 88.7 87.7 68.8
0.372 0.001 100 88.5 70.5
GDSS (Jo et al., 2022) 1.0 2.900 0.003 95.7 98.5 86.3
HI-Mol (Ours; 2%) 1.9 0.601 0.002 100 100 100
Appendix J Modification Algorithm
@@ Appendix I Details on QM9 Experiments
In Table 21, we report the full results of low-shot molecular property prediction experiments with averages and 95% confidence intervals. With randomly sampled low-shot molecules from the train split (used in our main experiments of Table 2), we generate $\times$3 number of valid molecules via generative models, e.g., we generate 96 molecules for 32-shot experiments. For the classifier, we utilize the 5-layer GIN (Xu et al., 2019a) from You et al. (2020), which is pre-trained with unlabeled molecules via self-supervised contrastive learning. We fine-tune this model for 100 epochs by introducing a linear projection head for each dataset. We use Adam optimizer with a learning rate of 0.0001 and no weight decay. The results are calculated based on the test ROC-AUC score of the epoch with the best validation ROC-AUC score. Specifically, we consider two scenarios: (1) training the classifier with only the low-shot molecules and (2) training the classifier with both the original low-shot molecules and the generated molecules via the molecular generative model. We report $\Delta$ROC-AUC score, calculated by the subtraction of the ROC-AUC score of (1) from (2).

### v2

@@ Abstract
Developing an effective molecular generation framework even with a limited number of molecules is often important for its practical deployment, e.g., drug discovery, since acquiring task-related molecular data requires expensive and time-consuming experimental costs. To tackle this issue, we introduce Hierarchical textual Inversion for Molecular generation (HI-Mol), a novel data-efficient molecular generation method. HI-Mol is inspired by the importance of hierarchical information, e.g., both coarse- and fine-grained features, in understanding the molecule distribution. We propose to use multi-level embeddings to reflect such hierarchical features based on the adoption of the recent textual inversion technique in the visual domain, which achieves data-efficient image generation. Compared to the conventional textual inversion method in the image domain using a single-level token embedding, our multi-level token embeddings allow the model to effectively learn the underlying low-shot molecule distribution. We then generate molecules based on the interpolation of the multi-level token embeddings. Extensive experiments demonstrate the superiority of HI-Mol with notable data-efficiency. For instance, on QM9, HI-Mol outperforms the prior state-of-the-art method with 50$\times$ less training data. We also show the effectiveness of molecules generated by HI-Mol in low-shot molecular property prediction. Code is available at https://github.com/Seojin-Kim/HI-Mol.
@@ 1 Introduction
Figure 1: Overview of HI-Mol framework. (1) Hierarchical textual inversion: we encode low-shot molecules into multi-level token embeddings. (2) Embedding interpolation-based sampling: we generate novel molecules using interpolation of low-level token embeddings.
@@
We learn this hierarchical embedding in the token space of the recent text-to-molecule model (Edwards et al., 2022). Specifically, we adopt textual inversion (Gal et al., 2022)—which learns low-shot image distribution by introducing a new single token within a text-to-image model. A key difference here is that the consideration of “multi-level” tokens (see Figure 1) are essential in learning the low-shot molecule distribution due to the complicatedly structured nature of molecules compared to images (see Table 1).
@@ 2 Related Work
Molecular generation. Most molecular generation methods fall into three categories. First, graph-based methods (Jo et al., 2022; Hoogeboom et al., 2022; Luo et al., 2022; Zhang et al., 2023; Vignac et al., 2023) formalize molecular generation as a graph generation problem by representing each molecule as an attributed graph. Next, fragment-based methods (Jin et al., 2018; Kong et al., 2022; Geng et al., 2023) define a dictionary of chemically meaningful fragments, e.g., functional groups. A molecule is represented as a tree structure of dictionary elements and the distribution of connected fragments is then modeled. Finally, string-based methods (Gómez-Bombarelli et al., 2016; Flam-Shepherd et al., 2022; Ahn et al., 2022) utilize the Simplified Molecular-Input Line-Entry System (SMILES, Weininger, 1988) to write molecules as strings and learn the distribution of molecules in this string space. Our method takes the string-based approach based on the recent large-scale text-to-molecule models that use the SMILES representation.
Hierarchical generation methods. Recent molecular generation methods introduce the notion of hierarchy in molecular generation (Jin et al., 2020; Zhu et al., 2023). Specifically, they incorporate the hierarchy within a single molecule, e.g., atom- and motif-level. However, they do not consider the hierarchy among molecules, e.g., dataset-, cluster-, and molecule-level, which is indeed crucial to understand the molecular dataset, i.e., target distribution (Alexander et al., 2011). To overcome this limitation, we carefully design our multi-level embeddings to reflect the hierarchy among the target molecules through our hierarchical tokens.
@@
In contrast to the recent advances of low-shot generation in the image domain, developing a low-shot (or data-efficient) molecular generation framework is relatively under-explored despite its importance in practical applications (Altae-Tran et al., 2017; Guo et al., 2022). Hence, our method tackles this problem by designing a molecule-specific inversion method using the recent large-scale text-to-molecule models. Specifically, we incorporate the concept of “hierarchy” of molecular structures (Alexander et al., 2011) into the textual inversion framework. Due to this unique motivation, our method effectively learns the molecule distribution with the shared concept in low-shot molecules with diverse molecular structures, while the applications of prior works, e.g., Guo et al. (2022), are limited to structurally similar low-shot molecules such as monomers and chain-extenders (see Table 7 for comparison).
@@ 3 HI-Mol: Hierarchical Textual Inversion for Molecular Generation
In Section 3.2, we provide descriptions of textual inversion to explain our method. In Section 3.1, we provide an overview of our problem and the main idea. In Section 3.3, we provide a component-wise description of our method.
@@ 4.2 Main Results
We also propose a simple algorithm to modify the generated invalid SMILES strings by correcting invalid patterns [^5] For example, we modify the invalid SMILES caused by the unclosed ring, e.g., $\mathtt{C1CCC}\rightarrow\mathtt{CCCCC}$. Please see Appendix I for the detailed algorithm. We mark in the Grammar column in Table 2 and 4 when modification is applied for evaluation. without a computational overhead. By applying this modification algorithm, we convert an invalid SMILES string to a valid one that represents a valid molecule, therefore, the Validity score becomes 100 in this case. In particular, the molecules from the modified SMILES further improve the overall metrics, e.g., FCD by 19.0 $\rightarrow$ 16.6 and 11.2 $\rightarrow$ 10.7 in the HIV and the BBBP dataset, respectively. This indicates that the modified SMILES indeed represent molecules from the desired low-shot molecule distribution and further highlights the superior quality of our generated molecules.
@@
For an extensive comparison with the baselines which show high Uniqueness and Novelty scores, e.g., GDSS (Jo et al., 2022), we perform an additional comparison after we adjust the Uniqueness and Novelty scores of our method to 100; this setup allows us to perform a fair comparison in the FCD score with these methods. Here, we adjust the sampling strategy slightly; we ignore the generated molecules which have an overlap with the training molecules and the already generated molecules. Even in this case, HI-Mol achieves an FCD of 0.601, which outperforms all these baselines. We provide detailed results and discussion in Appendix J.
@@
Extremely limited data regime. Since our model exploits the power of large molecular language models by designing a molecule-specialized textual inversion scheme, one can expect our model to be beneficial in extremely limited data regimes compared with prior methods. To verify this, we conduct an experiment using only subset of the HIV dataset and report its quantitative result in Table 7. Even with this situation, HI-Mol still outperforms prior state-of-the-art molecular generation methods, e.g., our method improves FCD as 39.2 $\to$ 34.8 when trained with 30 samples.
@@ 4.3 Analysis
Table 7: Results of molecular generation on subsets of the HIV dataset (Wu et al., 2018). We generate the same number of molecules as the number of the training samples. Due to the large training cost, we report the score of DEG (Guo et al., 2022) only for 30 samples.
30 DEG (Guo et al., 2022) Graph 03.3 39.2 0.105 100 100 100
STGG (Ahn et al., 2022) SMILES 00.0 41.5 0.110 100 67 100
CRNN (Segler et al., 2018) SMILES 00.0 40.0 0.121 80 71 100
HI-Mol (Ours) SMILES 08.3 34.8 0.103 80 75 100
150 STGG (Ahn et al., 2022) SMILES 01.3 28.2 0.054 100 90 100
CRNN (Segler et al., 2018) SMILES 01.3 30.1 0.063 50 84 100
HI-Mol (Ours) SMILES 08.3 22.1 0.038 64 91 100
500 STGG (Ahn et al., 2022) SMILES 01.3 22.8 0.041 100 74 100
CRNN (Segler et al., 2018) SMILES 02.7 30.0 0.064 51 100 100
HI-Mol (Ours) SMILES 10.3 20.8 0.020 63 91 100
Ablation on hierarchical textual inversion. To validate the effectiveness of each component in our HI-Mol framework, we compare the results where some components are excluded from the overall framework. Specifically, we compare the generation performance of the following setups: (1) not using the inversion technique; we train the text-to-molecule model with the molecule-description pairs, (2) using the shared token $[S^{\ast}]$ only, (3) using $[S^{\ast}]$ and the detail tokens $[D_{n}^{\ast}]$, (4) using all three types of tokens, and (5) applying the additional modification algorithm. Note that for (1) and (2), it is impossible to apply our interpolation-based sampling; instead, we use temperature sampling with temperature $\tau=2.0$. We provide this result in Table 6. First, we find that (1) the naïve training and (2) the inversion with a single shared token (Gal et al., 2022) do not show reasonable performance, i.e., they achieve only 0.4% Validity. In (3) and (4), introducing low-level tokens in the inversion framework significantly improves the generation quality by learning the low-level features in molecules. Finally, (5) the modification algorithm converts an invalid generated SMILES into a valid one that lies in our target distribution. We provide additional ablation results in Appendix G.
@@ 5 Conclusion
Acknowledgements
We thank all the anonymous reviewers for their insightful feedbacks and discussions. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST), No.RS-2021-II212068, Artificial Intelligence Innovation Hub) and Mogam institute for Biomedical Research and GC Biopharma.
@@ Appendix A Method Details
We utilize a recently introduced text-to-molecule model, MolT5-Large-Caption2Smiles (Edwards et al., 2022) in our HI-Mol framework. [^7] https://huggingface.co/laituan245/molt5-large-caption2smiles This model is constructed upon a text-to-text model, T5 (Raffel et al., 2020), and molecular information is injected by additional training with both unpaired SMILES (Weininger, 1988) string and caption-SMILES paired dataset. We update the token embeddings and the linear heads, while freezing other parameters. Our experiment is conducted for 1,000 epochs using a single NVIDIA GeForce RTX 3090 GPU with a batch size of 4. We use AdamW optimizer with $\epsilon=1.0\times 10^{-8}$ and let the learning rate $0.3$ with linear scheduler. We clip gradients with the maximum norm of 1.0. We update the assigned cluster $c_{n}$ of each molecule for the first 5 epochs following Eq. (3.3). For interpolation-based sampling, we choose a uniform distribution $p(\lambda)$, (i.e., $p(\lambda)\coloneqq\mathcal{U}(l,1-l)$), where $\lambda$ controls relative contributions of interpolated token embeddings. We set $l=0.0$ on the datasets in MoleculeNet benchmark (Wu et al., 2018), and $l=0.3$ on the QM9 dataset (Ramakrishnan et al., 2014).
@@ Appendix G Ablation Study
Table 12: Ablation of hierarchical textual inversion on the HIV dataset in the MoleculeNet (Wu et al., 2018) benchmark.
Inverted tokens Active. $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Valid. $\uparrow$ Unique. $\uparrow$ Novelty $\uparrow$
$[D^{\ast}]$ 5.4 21.7 0.026 100 88.8 100
$[S^{\ast}][I^{\ast}][D^{\ast}]$ 11.4 16.6 0.019 100 95.6 100
@@
Effect of hierarchical textual inversion. In Table 12, we show the importance of our hierarchical textual inversion. Specifically, we compare using a single hierarchy, i.e., detail tokens (using $[D^{\ast}]$), and our multi-level hierarchy (using $[S^{\ast}][I^{\ast}][D^{\ast}]$). The results show that our multi-level textual inversion strategy is highly useful to generate faithful molecules from the desired distribution.
@@ Appendix H Additional Experiments
Appendix I Modification Algorithm
@@
Appendix J Details on QM9 Experiments
[…]
2 GDSS (Jo et al., 2022) 22.953 0.455 99.8 1.2 72.2
STGG (Ahn et al., 2022) 0.715 0.002 100 88.0 42.1
HI-Mol (Ours) 0.434 0.001 90.7 75.8 73.5
HI-Mol (Ours) 0.430 0.001 100 76.1 75.6
5 GDSS (Jo et al., 2022) 17.013 0.066 97.2 25.5 44.2
STGG (Ahn et al., 2022) 0.665 0.001 100 95.8 63.0
HI-Mol (Ours) 0.412 0.001 89.4 85.8 70.4
HI-Mol (Ours) 0.410 0.001 100 86.4 72.4
10 GDSS (Jo et al., 2022) 17.170 0.067 98.0 22.8 36.6
STGG (Ahn et al., 2022) 0.603 0.002 100 99.4 63.5
HI-Mol (Ours) 0.400 0.002 87.6 87.6 71.2
HI-Mol (Ours) 0.398 0.001 100 88.3 73.2
20 GDSS (Jo et al., 2022) 7.345 0.025 94.2 82.3 67.6
STGG (Ahn et al., 2022) 0.599 0.001 100 99.4 64.3
HI-Mol (Ours) 0.384 0.001 86.7 87.8 70.0
HI-Mol (Ours) 0.383 0.001 100 88.7 71.8
50 GDSS (Jo et al., 2022) 3.564 0.008 96.0 96.6 80.1
STGG (Ahn et al., 2022) 0.592 0.001 100 99.2 70.6
HI-Mol (Ours) 0.372 0.001 88.7 87.7 68.8
HI-Mol (Ours) 0.372 0.001 100 88.5 70.5
GDSS (Jo et al., 2022) 1.0 2.900 0.003 95.7 98.5 86.3
HI-Mol (Ours; 2%) 1.9 0.601 0.002 100 100 100
@@
Table 23: Comparison with latent mixup (Wang et al., 2021) in the low-shot classification task. We report $\Delta$ROC-AUC averaged over 20 random seeds.
32-Shot HIV BBBP BACE
Latent mixup (Wang et al., 2021) 0.55 1.27 0.52
HI-Mol (Ours) 2.16 2.64 3.39
Low-shot (or few-shot) prediction tasks are one of the important applications for industrial deployments (Nam et al., 2023), and we have shown our HI-Mol’s capability to be beneficial to these tasks. In Table 22, we report the full results of low-shot molecular property prediction experiments with averages and 95% confidence intervals. With randomly sampled low-shot molecules from the train split (used in our main experiments of Table 2), we generate $\times$3 number of valid molecules via generative models, e.g., we generate 96 molecules for 32-shot experiments. For the classifier, we utilize the 5-layer GIN (Xu et al., 2019a) from You et al. (2020), which is pre-trained with unlabeled molecules via self-supervised contrastive learning. We fine-tune this model for 100 epochs by introducing a linear projection head for each dataset. We use Adam optimizer with a learning rate of 0.0001 and no weight decay. The results are calculated based on the test ROC-AUC score of the epoch with the best validation ROC-AUC score. Specifically, we consider two scenarios: (1) training the classifier with only the low-shot molecules and (2) training the classifier with both the original low-shot molecules and the generated molecules via the molecular generative model. We report $\Delta$ROC-AUC score, calculated by the subtraction of the ROC-AUC score of (1) from (2). In Table 23, we additionally compare with conventional latent mixup strategy (Wang et al., 2021). They directly use the interpolated latent embeddings (and corresponding interpolated labels) as inputs, which mostly do not become real data. However, we generate “new molecules” (rather than just latent embeddings) based on this embedding and use it as real data to train a classifier for a molecular prediction task. For the latent mixup, we train the classifier using given molecules and interpolated latent embeddings (and labels) using uniformly sampling coefficient $\lambda$ from a range of [0, 1] (which is the same with the choice of $\lambda$ in our method). As shown in the table, our method indeed shows a significantly better $\Delta$ROC-AUC compared to latent mixup.